OpenStudy (anonymous):

Prove: tan^2 θ•cos^2 θ + cos^2 θ = 1. You must show all work. any help please?

OpenStudy (anonymous):

Okay so this is what grade?

Nnesha (nnesha):

write tan^2 in terms of sin cosine tan = what ?

OpenStudy (anonymous):

this is 11th grade math. trigonometry in Algebra 2

OpenStudy (anonymous):

i have no clue

Nnesha (nnesha):

alright first you need to know the definition of sin /cos and tan 1/csc ,1/ sec ,1/cot are reciprocal of sin , cos and sin \[\tan =\frac{ 1 }{ \cot }\] or you can say tan =sin/cos same thing

Nnesha (nnesha):

\[\tan=\frac{ \sin }{ \cos } ~~~\cos= \frac{ 1 }{ \sec }~~~\sin=\frac{ 1 }{ \csc }\] for that question we just need tan=sin/cos

OpenStudy (anonymous):

oh, right... x/y

Nnesha (nnesha):

\[\huge\rm \color{ReD}{tan^2 \theta} • \cos^2 \theta+\cos^2 \theta=1\] we can rewrite tan^2 in terms of sin cosine \[\huge\rm \color{Red}{\frac{ \sin^2 \theta }{ \cos^2 \theta }}• \cos^2 \theta+\cos^2 \theta=1\]

OpenStudy (anonymous):

which would eliminate the cos^2

Nnesha (nnesha):

yes right \[\huge\rm \color{Red}{\frac{\sin^2 \theta }{\cancel{ \cos^2 \theta} }}• \cancel {\cos^2 \theta}+\cos^2 \theta=1\] left with \[\huge\rm\color{blue}{ sin^2 \theta +\cos^2} =1\] you should b familiar with that special identity

OpenStudy (anonymous):

okay, thanks

Nnesha (nnesha):

cos^2 theta **

Nnesha (nnesha):

sin^2 +cos^2 equal to what ??

imqwerty (imqwerty):

1

imqwerty (imqwerty):

hey m jst guessing ( ͡° ͜ʖ ͡° )